OpenStudy (anonymous):

what is the slope of the line that passes through the given points (2,12) and (6,11) ? * -1/4 *1/4 *4 *-4

OpenStudy (one098):

-4

OpenStudy (one098):

2-6/12-11=-4/1=-4

OpenStudy (anonymous):

thanks(:

OpenStudy (one098):

You're welcome, anytime!